OpenStudy (anonymous):

A 10 foot ladder is leaned against a building so that the distance from the bottom of the ladder to the building is 2 feet less than the distance from the top of the ladder to the ground. What is the distance from the bottom of the ladder to the building? Hint: you must use the Pythagorean theorem.You must construct an algebra equation that contains a variable.

OpenStudy (anonymous):

x^2+(x-2)^2 =100 then solve for x

OpenStudy (anonymous):

x = height of ladder from gound x-2 = dtsnce from ground x^2 + (x-2)^2 = 10^2 solve for x

OpenStudy (anonymous):

You have to solve for x-2, not x if those are your chosen lengths.

OpenStudy (anonymous):

solving will give x= 8

OpenStudy (anonymous):

right - after you have x then distance from ground = x -2

OpenStudy (anonymous):

as expanding the equation that will be shown as : x^2-2x-48=0 ===> x=8 or x=-6 ===> so x=8